a vehicle is travelling at 70km/h and a child ran onto the road 120m in front. can the driver avoid the child.


1
Expert's answer
2021-11-17T10:37:07-0500

"70\ km/h=19.4\ m/s"


"s=\frac{v^2-v_0^2}{2a}\to a=\frac{v^2-v_0^2}{2s}=\frac{0^2-19.4^2}{2\cdot 120}=-1.57\ (m/s^2)"


The driver will be able to avoid the child if he starts to brake with acceleration "|a|\geq 1.57\ (m/s^2)" .